OpenStudy (anonymous):

Square root of 8/50. Please explain the process in your answer.

jimthompson5910 (jim_thompson5910):

Hint: 8/50 = 4/25

OpenStudy (anonymous):

therefore 2/5 as the root of 4 is 2 and the root of 25 is 5

OpenStudy (anonymous):

Sorry but I'm still confused. The last time I took a math class was last summer.

jimthompson5910 (jim_thompson5910):

did you see how I got 8/50 = 4/25 ?

OpenStudy (anonymous):

Divide it by 2.

jimthompson5910 (jim_thompson5910):

yep, divide each part by 2 8/2 = 4 50/2 = 25 Now take the square root of 4 to get 2 then take the square root of 25 to get 5 So that's why the square root of 4/25 = 2/5

OpenStudy (anonymous):

Oh, so reduce then get the square root of that?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

its always a good idea to reduce fractions as much as possible no matter what you're doing in math